OpenStudy (anonymous):

simplify (3x^5 x -4x^-8)^-2

OpenStudy (anonymous):

Seems like you got an extra x in parentheses

OpenStudy (anonymous):

thats a multiplication sytmbol

OpenStudy (lalaly):

\[(3 \times -4 x^{(5-8)})^{-2}\]\[=(-12x^{-3})^{-2}\]\[=(\frac{-12}{x^3})^{-2}\]\[=(\frac{-x^3}{12})^2\]\[=\frac{x^6}{144}\]

OpenStudy (anonymous):

thanks

OpenStudy (lalaly):

=)